# suppSupplementary References

### v1

@@ 4.3 Surface Area Regularization
Furthermore, we can show that the coarea loss is too aggressive for our purpose of signed distance field reconstruction. If we directly minimize the coarea loss (the left hand side in Eq. 16) by applying the Euler-Lagrange equation, the minimizer would satisfy:
$1-\frac{\nabla^{2}u}{\|\nabla u\|}=0,$ (17)
which does not relate to the regularization parameter $\lambda$. Unfortunately, the partial differential equation system formed by the equation above and the eikonal equation does not have a solution in general. We can see this by plugging in $\|\nabla u\|=1$ to get a Poisson’s equation $\nabla^{2}u=1$. Under certain boundary condition, it has a unique solution of the gradient $\nabla u$, while the gradient is not guaranteed to have a unit norm. This makes the coarea loss unable to accurately represent a signed distance function. By contrast, our heat loss has bounds for its distance approximation and converges to the true distance as we increase the parameter $\lambda$ (Eq. 9).
@@ 4.4 Relation to PHASE [44]
While derived from very different mathematical principles and a different motivation, our method has a close relation to the PHASE model proposed by Lipman. PHASE is motivated by smoothly approximating an occupancy function $o(\bm{x})$, which outputs $1$ outside the object, $-1$ inside, and 00 on the surface.
In the supplementary material, we show that there is a close relation between our heat field $h$ and the smoothed occupancy function $o$. We further show that our theory and derivation have led to additional insights and significant differences in implementation and results. In particular, PHASE’s theory introduces a condition stating that the boundary loss weight $w_{b}$ should be small, whereas our theory emphasizes the need for a firm boundary condition (Eq. 4 and Fig. 5).
Furthermore, we found that optimizing occupancy field in the PHASE model can be numerically unstable when the regularized occupancy $o$ approaches $1$ or $-1$. When used together with the eikonal loss, it can lead to unbounded network weight derivatives, impeding the optimization. Optimizing directly for the signed distance function like in our method avoids this issue.
@@ 5 Experiments
We show experiments on both 2D and 3D datasets. We construct a 2D dataset containing 14 shapes, ranging from simple geometric shapes to complex nested strctures. We also evaluate on two 3D benchmarks: a subset of ShapeNet [47] and the Surface Reconstruction Benchmark (SRB) [48]. Additionally, we generate a dataset with high genus geometries from Mehta et al. [1]. In all of our experiments, we use only the position information of the point cloud.
@@
We use a network with 5 hidden layers and 128 hidden channels for our method in all experiments. The hyperparameters are consistent within each dataset. To sample the integrals of our loss, we apply importance sampling to accurately estimate the heat loss integral. Full training details can be found in the supplementary material.
@@
An overview of our 2D dataset is attached in the supplementary material. For 2D experiments, we simply use a linear network and geometric initialization, as in DiGS [3]. We compare our method with DiGS [3] and StEik [4] with the hyperparameters and training configuration provided in their released code, and report the surface reconstruction metrics and distance field metrics in Table 2. Our method outperforms theirs across all metrics with only 8,192 sample points, whereas both of the previous methods use 15,000 sample points based on their default settings.
@@ 5.2 Ablation Study on 2D Dataset
The last row corresponds to our HotSpot model, which uses a combination of boundary, eikonal, and heat losses. Our method reconstructs the correct topology for all 14 shapes, while others struggle with the harder ones. It outperforms across nearly all metrics, demonstrating superior reconstruction quality and distance query accuracy. Although HotSpot slightly lags behind SAL [49] in RMSE and MAE, it successfully captures the correct topology for complex shapes, while SAL fails to do so. A visualization is shown in Fig. 6. Full qualitative results and detailed experiment parameters are provided in the supplementary material. Models relying heavily on area losses [3, 4] often face a trade-off between preserving original details and eliminating excess boundaries, which distorts the distance field (Section 4.3). Moreover, methods without our heat loss fall in local optimum easily.
@@
$\mathcal{L}_{B}$ $\mathcal{L}_{E}$ $\mathcal{L}_{A}$ $\mathcal{L}_{D}$ $\mathcal{L}_{DD}$ $\mathcal{L}_{S}$ $\mathcal{L}_{H}$ IoU $d_{C}$ $d_{H}$ RMSE MAE SMAPE
@@
Table 2: Ablation study of different loss combinations and metrics. From top to bottom, the loss combinations are from: IGR [7], SAL [49], DiGS w/o area loss, DiGS [3], StEik w/o loss, StEik [4], and our HotSpot. From left to right, the losses are boundary loss, eikonal loss, area loss, divergence loss, directional divergence loss, SAL loss, heat loss.
@@
We evaluate our method on a processed subset [19, 16, 50] of ShapeNet [47], which provides surface point sampled on preprocessed watertight meshes for 260 shapes across 13 categories. To ensure a reasonable spatial density for passing the heat, we scale the point cloud so that 70% of the points lie within a sphere centered at the origin with a radius of 0.45. For a fair comparison, we transform our outcomes back into the coordinate system used in StEik [4] and DiGS [3]. Additionally, we implemented a scheduler to gradually increase the absorption parameter $\lambda$, which enhances the representation of level set details.
We present the quantitative results in Table 3 and some visual results in Fig. 8. Our HotSpot model outperforms current state-of-the-art models across all surface reconstruction metrics, demonstrating that heat can effectively connect points, enabling the neural network to interpolate between them and produce high-quality surfaces. We also achieve near-top accuracy in distance queries. While SAL [49] uses the distance to the closest point in their loss function, leading to slightly lower absolute error across the entire space, their performance drops significantly in relative error and near-surface metrics. This suggests that the distance to the closest point alone is not a reliable approximation near the surface. In contrast, our neural network completes surfaces between nearby points, and our heat loss propagates heat smoothly, preserving high-quality level sets.
Figure 7: Visual examples on high genus dataset. Our method uses significantly less iterations but successfully reconstructs the correct topology for all shapes, whereas generate extra boundaries.
Fig. 1 and Fig. 7 shows the reconstruction of high genus surfaces from Mehta et al. [1]. Existing methods struggle with the complex topology, while our method can accurately reconstruct both the surface and the signed distance function.
In the supplementary material, we further present evaluations on other high genus shapes from Mehta et al. and the SRB dataset.
@@
Table 3: Surface reconstruction metrics on ShapeNet [47], where $d_{C}$ and $d_{H}$ stand for the Chamfer and Hausdorff distances, RMSE, MAE, and SMAPE are distance field metrics, and RMSE0.1, MAE0.1, and SMAPE0.1 are the same distance field metrics but only for points within distance $0.1$ of the surface. See the supplementary material for detailed results.
@@ 5.4 Sphere Tracing
Fig. 9 shows a visualization of number of steps a sphere tracer [21] takes to render the shapes obtained by SAL, DiGS, StEik, and our method. Our method leads to the least amount of iterations on average, since it generates smooth and accurate signed distance functions, which are crucial for rendering high-quality surfaces. More results can be found in the supplementary material.
Figure 9: Visualizations and statistics of sphere tracing [21] iteration counts required to locate the surface. In the visual results brighter pixels indicating higher iteration counts. The histograms show the distribution of iterations for all pixels across 10 different camera poses. HotSpot finds the surface in fewer steps.
@@ 6 Discussion
Determining absorption coefficient $\lambda$. The absorption coefficient $\lambda$ has the units of $[\text{L}]^{-1}$, where L denotes a length unit. Rescaling the space is equivalent to rescaling $\lambda$ in the equation. While a larger $\lambda$ can improve accuracy, it introduces two challenges. First, our heat loss includes the factor $e^{-2\lambda|u|}$, which may become smaller than the precision limit of floating-point numbers, leading to round-off errors. Second, even if round-off errors are ignored, a large $\lambda|u|$ may reduce the optimizer’s ability to accurately shape the field. However, even with this, our loss will still push $u$ away from 00, preventing the introduction of extra surfaces where $\lambda|u|$ should be large. We compensate by adding the eikonal loss together with our heat loss. When $\lambda|u|$ is large, the eikonal loss becomes the dominant term. Our scheduler for $\lambda$ also effectively shapes the distant regions. Spatially-adaptive parameters can potentially further improve the results.
Necessity of firm boundary condition. The heat diffusion of the screened Poisson equation is based on a well established boundary condition, and we enforce the boundary condition through the boundary loss over a discrete set of points, and rely on the spectral bias of neural networks [51] to interpolate them. When the input point set is sparse and the absorption coefficient $\lambda$ is high, an overly strong heat loss ($w_{b}$ being too small) can tear the boundary, causing the signed distance function to degrade into an unsigned distance one like Fig. 5. Thus, our theory suggests that a key insight is to set a high boundary weight $w_{b}$, while adjusting the absorption coefficient $\lambda$ to match the input point cloud density or by scaling the point cloud itself. Our experiments show that with proper parameter setting and rescaling the collapse can be avoided. Future research is required for very sparse boundary, and application to inverse rendering.
@@ 7 Conclusion
We propose a new model for neural signed distance function optimization based on screened Poisson equation. We analyze our loss theoretically and show that it can converge to the true distance, is stable both to a small perturbation and in the temporal dynamics, and penalizes large surface area. Our experiments show that our method reconstructs both better surfaces and better distance approximations compared to many existing methods, especially on complex and high-genus shapes.
@@ Appendix A Proofs and Derivations
A.1 Proposition 1 proof.
Here we analyze the 2D case, but the following derivation naturally generalizes to 3D. We denote $\frac{\partial u}{\partial x}$ as $p_{1}$, $\frac{\partial u}{\partial y}$ as $p_{2}$, and let $n^{2}(x,y)=p_{1}^{2}+p_{2}^{2}$. Following the derivation from Kulyabov et al. [52], we can obtain the characteristics equations as follows. For any $a\neq 0$ if any field $u$ satisfies the eikonal equation,
@@
We assume $s\in(0,M)$ ($M>0$ can be infinity) is a differentiable domain which makes every derivative exist and all of these equations hold. Then we prove such a parametric curve is a ray or a segment.
@@
Then the disturbed $h^{\prime}=h_{e}+h$ still satisfies the screened Poisson equation in $\mathbb{R}^{d}\setminus(B\cup\Gamma)$. Given the spherical condition, we can analytically solve this equation by transforming it to an ordinary differential equation. Knowing that $\nabla^{2}h(r)=r^{-2}\cdot d(r^{2}\frac{dh}{dr})/dr$, we have:
@@
$h(r)=\frac{e^{-\lambda\epsilon}}{K_{0}(\lambda\epsilon)}K_{0}(\lambda r)$ (28)
It should be noted that $K_{0}$ converges to 00 when $\lambda r$ goes infinity. More precisely, $K_{0}(\lambda r)$ and $e^{-\lambda r}/\sqrt{\lambda r}$ are infinitesimals of the same order [53]. Related property then becomes similar to 3D case.
@@ Appendix B Experiments
B.1 2D Dataset
@@
Figure 10: This figure illustrates all the 14 vector shapes in our 2D dataset. Each visualization is within the range ${[-1.2,1.2]}^{2}$. Spanning a spectrum from simple to complex, the dataset encompasses various topologies, smooth and irregular boundaries, as well as configurations with single and multiple objects. Our HotSpot model accurately reconstructs all the curves while preserving their correct topologies.
@@
All other models make some errors in topology. Even when their topologies are correct, details like the Target and House shapes in the first two columns are missing. When using a relatively smaller learning rate in the ablation study for DiGS and StEik, the outputs become overly flat, despite maintaining the same ratio among the loss weights from their paper. This can be interpreted as a limitation of their derivative-based losses, which, as a corollary of the eikonal equation, only serve as a necessary condition for the equation, encouraging the condition $|\nabla u|=c$, where $c$ can be any constant. With a small learning rate, their losses trap the outputs to $|\nabla u|=0$.
@@
From Fig. 3 and Fig. 12, we observe the influence from different $\lambda$. When $\lambda$ is very small, heat from the boundaries diffuses to distant regions, causing $h\approx 1$ almost everywhere in the test region. As a result, $u\approx 0$ across the domain, leading to an overly flat signed distance function with many extra boundaries. As $\lambda$ increases, the outputs become more regular. Notably, even without the eikonal loss, setting $\lambda=10$ yields outputs that are more regular than several baselines in Fig. 11. However, as $\lambda$ continues to increase, the factor $e^{-\lambda|u|}$ in the loss computation diminishes rapidly, especially where $|u|$ is large initially or grows during training. This makes optimization without the eikonal loss increasingly challenging and less effective. For instance, in the $\lambda=50$ and $\lambda=100$ subfigures with $w_{e}=0$, the values in the upper-right region remain greater than $1.0$ even after 200k iterations. Although our neural network provides some output values in these regions, they are significantly larger than the ground truth.
@@
Table 4: Comparison of 2D dataset reconstruction results.
@@
B.2 ShapeNet
Figure 13: Additional results from the ShapeNet dataset. We present the bench, where none of the methods achieve a satisfactory surface reconstruction, and the airplane, where all methods successfully capture detailed structures.
We show full metrics for the ShapeNet dataset in Tables 6, 7, and 8. We compare our method with the state-of-the-art methods, including SAL [2], SIREN without normalization [31], DiGS [3], and StEik [4]. Our method outperforms all other methods in terms of surface reconstruction metrics, including IoU, Chamfer distance, and Hausdorff distance. In terms of distance query metrics, our method achieves the best performance in terms of RMSE, MAE, and SMAPE. In the near-surface region, our method also outperforms other methods in terms of RMSE, MAE, and SMAPE.

### v2

@@ 4.3 Surface Area Regularization
Furthermore, we can show that the coarea loss is too aggressive for our purpose of signed distance field reconstruction. If we directly minimize the coarea loss (the left hand side in Eq. 16), according to Euler-Lagrange equation, the minimizer would satisfy:
$1-\nabla^{2}u=0,$ (17)
which does not relate to the regularization parameter $\lambda$. Unfortunately, this partial differential equation system formed by the coarea loss above has a unique solution under Dirichlet boundary condition, and the gradient of its solution does not have a unit norm. This makes the coarea loss unable to accurately shape a signed distance function. By contrast, our heat loss has bounds for its distance approximation and converges to the true distance as we increase the parameter $\lambda$ (Eq. 9). Another simple way to verify our convergence and the compatibility with the eikonal equation is to rewrite Eq. 4 in terms of $u$ following Eq. 6:
${||\nabla u||}^{2}-1-\frac{\text{sgn(u)}}{\lambda}\nabla^{2}u=0.$ (18)
@@ 4.4 Relation to PHASE [49]
While derived from different mathematical principles and motivations, our method is closely related to the PHASE model proposed by Lipman. It is motivated by smoothly approximating an occupancy function $o(\bm{x})$, which is defined as $1$ outside, $-1$ inside, and 00 on the surface.
In the supplement, we show that there is a close relation between our heat field $h$ and the smoothed occupancy function $o$. We also show that our theory and derivation provide additional insights and significant differences in implementation and results. In particular, PHASE’s theory introduces a condition stating that the boundary loss weight $w_{b}$ should be small, whereas our theory emphasizes the need for a firm boundary condition (Eq. 4 and Fig. 6).
We further found that optimizing occupancy field in the PHASE model, when used together with the eikonal loss, can lead to unbounded network weight derivatives, impeding the optimization. Our method avoids this issue.
@@ 5 Experiments
We experiment on both 2D and 3D datasets. We construct a 2D dataset containing 14 shapes, ranging from simple geometric shapes to complex nested structures. We also evaluate on two 3D benchmarks: a subset of ShapeNet [55] and the Surface Reconstruction Benchmark (SRB) [56]. Additionally, we generate a dataset with high-genus geometries from Mehta et al. [1]. In all of our experiments, we use only the position information of the point cloud.
@@
We use a network with 5 hidden layers and 128 hidden channels for our method in all experiments. The hyperparameters are kept consistent within each dataset. We additionally apply importance sampling to our heat loss integral. Full training details can be found in the supplement.
@@
An overview of our 2D dataset is attached in the supplementary material. For 2D experiments, we simply use a linear network and geometric initialization, as in DiGS [3]. We compare our method with DiGS [3] and StEik [4] with the hyperparameters and training configuration provided in their released code, and report the surface reconstruction metrics and distance field metrics in Table 2. Our method outperforms theirs across all metrics with only 8,192 sample points, whereas both DiGS [3] and StEik [4] use 15,000 sample points based on their default settings.
@@ 5.2 Ablation Study on 2D Dataset
The last row corresponds to our HotSpot model, combining boundary, eikonal, and heat losses. Our method reconstructs the correct topology for all 14 shapes, while others struggle with the harder ones. It outperforms across nearly all metrics, demonstrating superior reconstruction quality and distance query accuracy. Although HotSpot slightly trails SAL [2] in RMSE and MAE, it captures the complex topology where SAL fails. A visualization is shown in Fig. 7. Full qualitative results and detailed experiment parameters are provided in the supplementary material. Models relying heavily on area losses [3, 4] often face a trade-off between preserving original details and eliminating redundant boundaries, which distorts the distance field (Section 4.3). Moreover, methods without our heat loss fall in local optimum easily.
@@
$\mathcal{L}_{B}$ $\mathcal{L}_{E}$ $\mathcal{L}_{A}$ $\mathcal{L}_{D}$ $\mathcal{L}_{DD}$ $\mathcal{L}_{S}$ $\mathcal{L}_{H}$ IoU $\uparrow$ $d_{C}$downarrow $d_{H}$downarrow RMSE $\downarrow$ MAE $\downarrow$ SMAPE $\downarrow$
@@
Table 2: Ablation study of different loss combinations and metrics. From top to bottom, the loss combinations are from: IGR [10], SAL [2], DiGS w/o area loss, DiGS [3], StEik w/o loss, StEik [4], and our HotSpot. From left to right, the losses are boundary loss, eikonal loss, area loss, divergence loss, directional divergence loss, SAL loss, and heat loss. Bold and underlined data: optimal; bold only: suboptimal. Same below.
@@
We evaluate our method on a processed subset [19, 16, 57] of ShapeNet [55], which provides surface points sampled on preprocessed watertight meshes for 260 shapes across 13 categories. To ensure a reasonable spatial density for passing the heat, we scale the point cloud so that 70% of the points lie within a sphere centered at the origin with a radius of 0.45. For a fair comparison, we transform our outcomes back into the coordinate system used in StEik [4] and DiGS [3]. Additionally, we implement a scheduler to gradually increase the absorption parameter $\lambda$, which enhances the representation of level set details.
We present the quantitative results in Table 3 and visual results in Fig. 9. More baselines [51, 46] are compared in the supplement. Our HotSpot model outperforms current state-of-the-art models across all surface reconstruction metrics, and achieves near-top accuracy in distance queries. While SAL [2] uses the closest point distance in their loss function, leading to slightly lower total absolute error, their performance drops significantly in relative error and near-surface metrics. This suggests that the closest point distance alone is unreliable near the surface. In contrast, as mentioned at the end of Section 3, after the interpolation driven by spectral bias [58], our loss remains effective in faithfully capturing the distance to the interpolated surface.
Figure 8: Visual examples on high genus dataset. Our method uses significantly fewer iterations but reconstructs the topology correctly for all shapes, whereas others generate extra boundaries.
Fig. 1 and Fig. 8 show high-genus surface reconstruction with shapes from Mehta et al. [1]. Existing methods struggle with the complex topology, while our method accurately reconstructs both the surface and the level set. In the supplement, we present evaluations on other high-genus shapes from Mehta et al. and the SRB dataset.
@@
Table 3: Surface reconstruction metrics on ShapeNet [55], where $d_{C}$ and $d_{H}$ stand for the Chamfer and Hausdorff distances, RMSE, MAE, and SMAPE are distance field metrics, and RMSE0.1, MAE0.1, and SMAPE0.1 are the same distance field metrics but only for points within distance $0.1$ of the surface. See the supplementary material for detailed results with more baselines.
@@ 5.4 Sphere Tracing
Fig. 10 visualizes the number of steps a sphere tracer [21] takes to render the shapes obtained by SAL, DiGS, StEik, and our method. Our method has the fewest iterations on average, since it provides the most accurate level sets. More results can be found in the supplement.
Figure 10: Visualizations and statistics of sphere tracing [21] iteration counts required to locate the surface. In the visual results brighter pixels indicate higher iteration counts. The histograms show the distribution of iterations for all pixels across 10 different camera poses. HotSpot finds the surface in fewer steps.
@@ 6 Discussion
Efficiency. Our method uses a loss function of only the neural network output and its first-order derivatives, and thus can be more efficient than previous methods which involve second-order derivatives. We report the runtime of our method and StEik [4] for training a model from ShapeNet [55] over 100 iterations using a single NVIDIA A10 GPU with varying network sizes, while keeping all other settings identical, as shown in Tab. 4.
Determining absorption coefficient $\lambda$. The absorption coefficient $\lambda$ has the units of $[\text{L}]^{-1}$, where L denotes a length unit. Rescaling the space is equivalent to rescaling $\lambda$. While a larger $\lambda$ can improve accuracy, it introduces two challenges: (1) our heat loss includes the factor $e^{-2\lambda|u|}$, which may become smaller than the precision limit of floating-point numbers, leading to round-off errors; (2) even if round-off errors are ignored, a large $\lambda|u|$ may reduce the optimizer’s ability to effectively shape the field. Nevertheless, our loss still pushes $u$ away from 00, preventing unwanted surfaces where $\lambda|u|$ should be large. We compensate by combining eikonal loss with heat losses. When $\lambda|u|$ is large, the eikonal loss becomes dominant. Our $\lambda$ scheduler also effectively shapes the distant regions. Spatially-adaptive parameters may also improve the results.
Structure $5\times 128$ $5\times 256$ $8\times 128$ $8\times 256$
StEik [4] 5.65s 10.62s 8.78s 17.00s
Ours 4.49s 8.43s 6.79s 12.86s
Table 4: Runtime for 100 iterations with different network sizes.
Necessity of firm boundary condition. The heat diffusion of the screened Poisson equation is based on a well-established boundary condition. We enforce the boundary condition through the boundary loss over discrete points, with neural networks interpolating through spectral bias [58]. When the input points are sparse and the absorption coefficient $\lambda$ is high, an overly strong heat loss ($w_{b}$ being too small) can tear the boundary, causing the signed distance to collapse into an unsigned one like Fig. 6. Thus, our theory highlights the importance of a high boundary weight $w_{b}$, while absorption $\lambda$ is adjusted to match the input point density or by scaling the point cloud itself. Our experiments show that proper parameter setting and rescaling prevent the collapse. Future research is required for very sparse boundaries, and applications to inverse rendering.
@@ 7 Conclusion
We propose a new model for neural signed distance function optimization based on the screened Poisson equation. We analyze our loss theoretically and show that it is an asymptotically sufficient condition to the true distance, is stable both to a small perturbation and in the temporal dynamics, and penalizes large surface area. Our experiments show that we reconstruct both better surfaces and better distance approximations compared to many existing methods, especially on complex and high-genus shapes.
Acknowledgements.
This work was supported in part by NSF grants 2127544, 2238839, 2100237, 2120019, and gifts from Adobe and Google. Additionally, we would like to thank Bing Xu, Zilu Li, Yash Belhe, Ishit Mehta, and Jianan Xiao for their suggestions, discussions, proofreading, and providing reference materials.
@@ Appendix A Proofs and Derivations
A.1 Proof of Proposition 1.
Here we analyze the 2D case, but the following derivation naturally generalizes to 3D. We denote $\frac{\partial u}{\partial x}$ as $p_{1}$, $\frac{\partial u}{\partial y}$ as $p_{2}$, and let $n^{2}(x,y)=p_{1}^{2}+p_{2}^{2}$. Following the derivation from Kulyabov et al. [59], we can obtain the characteristic equations as follows. For any $a\neq 0$ if any field $u$ satisfies the eikonal equation,
@@
We assume $s\in(0,M)$ ($M>0$ can be infinity) is a differentiable domain such that every derivative exists and all of these equations hold. Then we prove such a parametric curve is a ray or a segment.
@@
Then the perturbed $h^{\prime}=h_{e}+h$ still satisfies the screened Poisson equation in $\mathbb{R}^{d}\setminus(B\cup\Gamma)$. Given the spherical condition, we can analytically solve this equation by transforming it to an ordinary differential equation. Knowing that $\nabla^{2}h(r)=r^{-2}\cdot d(r^{2}\frac{dh}{dr})/dr$, we have:
@@
$h(r)=\frac{h_{0e}}{K_{0}(\lambda\epsilon)}K_{0}(\lambda r)$ (29)
It should be noted that $K_{0}$ converges to 00 when $\lambda r$ goes to infinity. More rigorously, $K_{0}(\lambda r)$ and $e^{-\lambda r}/\sqrt{\lambda r}$ are infinitesimals of the same order [60]. Related property then becomes similar to the 3D case.
@@ Appendix B Experiments
B.1 1D Verification
We can easily identify which loss function is not a sufficient constraint by observing its input. As discussed in the main text, if a loss function depends solely on the first-order derivative of the SDF or any higher-order derivatives, it is inherently insufficient as a constraint and cannot exclude other non-SDF solutions.
Our experiments also demonstrate that implementations optimizing SDFs with the closest point [2, 51, 52] may struggle when making queries near the interpolated surface. As mentioned at the end of Section 3, after the neural network interpolates these points due to spectral bias [58], our loss remains effective in faithfully capturing the distance to the interpolated surface.
For other cases, we can easily test whether a constraint is insufficient with simple 1D verifications.
We found one possible candidate to be a sufficient constraint from Marschner et al. [5] with following loss:
[…]
The loss is designed to optimize constructive solid geometries and formulated using both neural network outputs and their gradients, ensuring that when taking a step in the direction indicated by the SDF, the new position should have a distance value close to zero.
Figure 11: Marschner et al. [5] proposed a CP loss which is a possible candidate to be a sufficient constraint. However, our 1D experiment illustrates that it is incapable of converging to the actual signed distance function (dashed line), even when the output minimizes this loss almost everywhere. The $x$ axis is the domain and the $y$ axis shows the output of the implicit function. The middle two rows display the intermediate states of the optimization, while the bottom row presents the final results.
However, our experiment in Fig. 11 reveals that it cannot exclude non-SDF solutions like $u(\bm{x})\equiv 0$, confirming that it is not a sufficient constraint.
B.2 2D Dataset
@@
Figure 12: This figure illustrates all 14 vector shapes in our 2D dataset. Each visualization lies within the range ${[-1.2,1.2]}^{2}$. Spanning a spectrum from simple to complex, the dataset encompasses various topologies, smooth and irregular boundaries, as well as configurations with single and multiple objects. Our HotSpot model accurately reconstructs all curves while preserving their correct topologies.
@@
All other models make some errors in topology. Even when their topologies are correct, details like the Target and House shapes in the first two columns are missing. When using a relatively smaller learning rate in the ablation study for DiGS and StEik, the outputs become overly flat, despite maintaining the same ratio among the loss weights from their paper. This can be interpreted as a limitation of their derivative-based losses, which, as a corollary of the eikonal equation, only serve as a necessary condition for the equation, encouraging the condition $|\nabla u|=c$, where $c$ can be any constant. With a small learning rate, their losses trap the outputs at $|\nabla u|=0$.
@@
From Fig. 4 and Fig. 14, we observe the influence of different $\lambda$. When $\lambda$ is very small, heat from the boundaries diffuses to distant regions, causing $h\approx 1$ almost everywhere in the test region. As a result, $u\approx 0$ across the domain, leading to an overly flat signed distance function with many extra boundaries. As $\lambda$ increases, the outputs become more regular. Notably, even without the eikonal loss, setting $\lambda=10$ yields outputs that are more regular than several baselines in Fig. 13. However, as $\lambda$ continues to increase, the factor $e^{-\lambda|u|}$ in the loss computation diminishes rapidly, especially where $|u|$ is large initially or grows during training. This makes optimization without the eikonal loss increasingly challenging and less effective. For instance, in the $\lambda=50$ and $\lambda=100$ subfigures with $w_{e}=0$, the values in the upper-right region remain greater than $1.0$ even after 200k iterations. Although our neural network provides some output values in these regions, they are significantly larger than the ground truth.
@@
Table 5: Comparison of 2D dataset reconstruction metrics.
@@
B.3 ShapeNet
Figure 15: Additional results from the ShapeNet dataset [55]. We present the bench, where none of the methods achieves a satisfactory surface reconstruction, and the airplane, where all methods successfully capture detailed structures.
We show full metrics for the ShapeNet dataset in Tables 7, 8, and 9. We compare our method with the state-of-the-art methods, including SAL [2], SIREN without normalization [34], Neural-Singular-Hessian [45], Neural-Pull [51], DiGS [3], and StEik [4]. Our method outperforms all other methods in terms of surface reconstruction metrics, including IoU, Chamfer distance, and Hausdorff distance. In terms of distance query metrics, our method achieves near-top performance across RMSE, MAE, and SMAPE.
Notably, in near-surface regions, our approach outperforms all other methods, including SAL [2] and Neural-Singular-Hessian [45], across these metrics. As discussed in the main text, while approximating the SDF using the closest point information may be reasonably accurate for distant regions, it proves inadequate for near-surface regions. This limitation is particularly critical for sphere tracing—one of the most important applications of SDF—as it heavily relies on accurate field values in near-surface regions due to the high density of queries in these areas.
In contrast, as we introduced at the end of Section 3, after the neural network interpolates these points due to spectral bias [58], our loss remains effective in faithfully capturing the distance to the interpolated surface. Our experiments on sphere tracing further substantiate this claim, demonstrating the infeasibility of approximating the SDF using only the closest point information and highlighting the superiority of our proposed model.